OpenStudy (tanaciious):

cot^2x(cot^2x+1)=csc^2x/tan^2x

jhonyy9 (jhonyy9):

do you know csc x = ? how can you rewriting it in form of a fraction

jhonyy9 (jhonyy9):

the right side you can rewriting in this form csc^2 x *1/tan^2 x but you know that 1/tan x = cot x so in this way using this the above wrote 1/tan^2 x = cot^2 x so the right side will be csc^2 x *cot^2 x yes ? do you understand it ?

jhonyy9 (jhonyy9):

@Nnesha your opinion about above wrote please ? like a first step idea

OpenStudy (sshayer):

\[\csc ^2x-\cot ^2x=1,\csc ^2x=\cot ^2x+1\] \[\cot ^2x \left( \cot ^2x+1 \right)=\cot ^2x \csc ^2x=\frac{ \csc ^2x }{ \tan ^2x }\]